Is a triangle with sides of 2m,3m,and 4m a right triangle?

Answer by mathesther05(3)

a^2 +b^2 = c^2
2^2 + 3^2 = 4^2
4 + 9 = 16
13 < 16 when has to be 16 = 16 to have a right angle
therefore - this could not be a right angle triangle